OpenStudy (anonymous):

Find the sum of the arithmetic sequence. (1 point) -10, -7, -4, -1, 2, 5, 8

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (welshfella):

sum = (n/2) (a1 + l)

OpenStudy (anonymous):

i don't get what that equations mean lol

OpenStudy (welshfella):

where n = number of terms, a1 = first term and l = last term

OpenStudy (welshfella):

here n = 7 , a1 = -10 and l = 8

OpenStudy (welshfella):

just plug those values into the formula and work it out

OpenStudy (welshfella):

= (7/2) * (-10+8) = (7/2) * -2 = ?

OpenStudy (anonymous):

or, in this short case, you can add them one by one, may be a little bit faster ;)

OpenStudy (welshfella):

yes true - but its a good idea to remember the formula as well

OpenStudy (anonymous):

okay i have another question

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

?

Nnesha (nnesha):

for geometric sequence there are two formulas to find sum it depends on ratio if ratio is less than then you have to apply infinite geometric formula \[\large\rm \left| r \right| <1 \] \[\huge\rm s_n = \frac{ a_1 }{ 1-r }\] and if ratio is greater than one then use finite geo. \[\huge\rm s_n = a_1 ( \frac{ 1 -r^n }{ 1-r })\]

OpenStudy (anonymous):

its going smaller does that matter?

Nnesha (nnesha):

what is ratio ?? r= ?? do you know how to find ratio ?

OpenStudy (anonymous):

no

Nnesha (nnesha):

to find ratio divide next term by one before term \[\huge\rm r = \frac{ a_2 }{ a_1 } , \frac{ a_4 }{ a_3 }\]

OpenStudy (anonymous):

so is the answer 1/768?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

how did you get that ?

OpenStudy (anonymous):

i added all of them

Nnesha (nnesha):

r is less than one ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so apply infinite geo. formula

OpenStudy (anonymous):

so is my answer right?